# Pivoting technique for the circle homeomorphism group

### v1

@@ Abstract.
This is a translation of part of [Cho25]. We adapt Gouëzel’s pivoting technique to the circle homeomorphism group. As an application, we give different proofs of Gilabert Vio’s probabilistic Tits alternative and Malicet’s exponential synchronization.
@@ 1. Introduction
(Zn and Zn comprise a ping-pong $\operatorname{\mathbb{P}}\left(\textrm{$Z_{n}$ and $Z_{n}^{\prime}$ comprise aping-pong pair}\right)\geq 1-q^{n}$
@@
[…]
@@
(1.1) (Zn and Zn comprise a ping-pong $\operatorname{\mathbb{P}}\left(\textrm{$Z_{n}$ and $Z_{n}^{\prime}$ comprise aping-pong pair}\right)\geq 1-\frac{1}{\kappa}e^{-\kappa n}$
@@
In the first case, $f_{i}$’s simultaneously preserve a metric on $S^{1}$ and distinct points are kept distant. In the second case, a global synchronization cannot be expected but a local synchronization can happen. In the third case, synchronization happens almost surely. This was promoted into exponential synchronizing proven by Dominique Malicet. Note that the minimality assumption is lifted.
@@
We strengthen this result by providing an exponential bound on the probability:
@@
(1.2) $\operatorname{\mathbb{P}}\left(\operatorname{\omega}:\begin{array}[]{c}\textrm{there exists an interval $I_{x,\operatorname{\omega}}$ containing $x$ such that}\\ \textrm{ $\operatorname{\operatorname{diam}}\big{(}Z_{k}(\operatorname{\omega})(I_{x,\operatorname{\omega}})\big{)}\leq q^{k}$ for each $k\geq n$}\end{array}\right)\geq 1-\frac{1}{\kappa}e^{-\kappa n}$
@@
(1.3) $\operatorname{\mathbb{P}}\left(\operatorname{\omega}:\begin{array}[]{c}\textrm{there exists an interval $I_{x,\operatorname{\omega}}$ containing $x$ such that}\\ \textrm{$\operatorname{\operatorname{diam}}(I_{x,\operatorname{\omega}})\geq 1-q^{n}$ and $\operatorname{\operatorname{diam}}\big{(}Z_{k}(\operatorname{\omega})(I_{x,\operatorname{\omega}})\big{)}\leq q^{k}$ for each $k\geq n$}\end{array}\right)\geq 1-\frac{1}{\kappa}e^{-\kappa n}$
@@
$\operatorname{\mathbb{P}}\left(\operatorname{\omega}:\begin{array}[]{c}\textrm{there exists $\xi=\xi(\operatorname{\omega},n)\in\partial X$ such that $\big{(}Z_{k}(\operatorname{\omega})\xi^{\prime}\big{|}Z_{k}(\operatorname{\omega})o\big{)}_{o}>\kappa k$}\\ \textrm{ for every $k\geq n$ and for every $\xi^{\prime}\in\partial X$ such that $\big{(}\xi^{\prime}\big{|}\xi\big{)}_{o}<\kappa n$}\end{array}\right)\geq 1-\frac{1}{\kappa}e^{-\kappa n}$
@@ Acknowledgments
The author thanks Hyungryul Baik and Sang-hyun Kim for their valuable comments on the Korean version of this paper. The author thanks Dominque Malicet and Martín Gilabert Vio for explaining the basics of circle homeomorphisms and clarifying the author’s many confusions. Malicet and Gilabert Vio encouraged the author to consider non-degenerate random walks on subsemigroups (which are not necessarily subgroups) of $\operatorname{Homeo}(S^{1})$, which led to the current statements. The author is grateful to their suggestions. The author also thanks Gilabert Vio for many corrections and suggestions for the first draft of this paper.
The author was partially supported by Mid-Career Researcher Program (RS-2023-00278510) through the National Research Foundation funded by the government of Korea. The main part of the paper was written while the author was visiting the HCMC of KIAS.
@@ 2. Preliminaries
For distinct points $x,y\in S^{1}$, let $\tilde{x}$ and $\tilde{y}$ be lifts on $\mathbb{R}$ of $x$ and $y$, respectively, such that $\tilde{x}<\tilde{y}<\tilde{x}+1$. We define the open interval $(x,y)$ in $S^{1}$ by the image of $(\tilde{x},\tilde{y})\subseteq\mathbb{R}$ on $S^{1}$ by the quotient map. Equivalently,
@@
Given a subset $A\subseteq S^{1}$, we denote by $\zeta(A)$ the number of connected components of $A$ (possibly +$\infty$+ ∞). Hence, $\zeta(A)=1$ precisely when $A$ is connected, i.e., is a point, interval or $S^{1}$.
@@ 3. Weak Tits alternative for semigroups
    $G$ acts on $S^{1}$ proximally and every $G$-orbit is infinite, i.e., #$\#Gx=+\infty$ for each $x\in S^{1}$.
@@
(3.2) $N(A;I):=\sup\left\{\#S:\begin{array}[]{c}\textrm{$S\subseteq G$, $gI\subseteq A$ for each $g\in S$},\\ \textrm{ $gI\cap g^{\prime}I=\emptyset$ for distinct elements $g,g^{\prime}\in S$}\end{array}\right\}.$
@@
$\left(\frac{N(A;I_{n})}{N(S;I_{n})}\right)_{n>0}=\left(\frac{N(A;I_{1})}{N(S;I_{1})},\,\frac{N(A;I_{2})}{N(S;I_{2})},\,\frac{N(A;I_{3})}{N(S;I_{3})},\ldots\right)$
is well-defined thanks to Claim 3.6 and is bounded between 0 and 1. Hence, up to replacing $(I_{n})_{n>0}$ with its subsequence, $(N(A;I_{n})/N(S;I_{n}))$ converges. Since $\mathcal{E}$ is countable, we can subsequently take convergent subsequences for each $A\in\mathcal{E}$. As a result, for a suitable interval neighborhood basis $(I_{n})_{n>0}$ of $x$, we can guarantee that
$\lim_{n\rightarrow+\infty}\frac{N(A;I_{n})}{N(S;I_{n})}=:\mu(A)\,\,\textrm{exists for each $A\in\mathcal{E}$.}$
@@
Meanwhile, we observed in Claim 3.5 that $N(S;I_{n})$ grows indefinitely as $n$ tends to infinity. Hence, we conclude
$00\leq\big{|}\mu(A_{1})+\mu(A_{2})-\mu(A)\big{|}=\lim_{n\rightarrow+\infty}\left|\frac{N(A_{1};I_{n})+N(A_{2};I_{n})-N(A;I_{n})}{N(S;I_{n})}\right|\leq\lim_{n\rightarrow+\infty}\left|\frac{2}{N(S;I_{n})}\right|=$ 0 .
@@
Each $A_{i}$ is of the form $(a_{i},b_{i}]$. We now take $c_{i}\in S^{1}$ such that $[b_{i},c_{i}]$ has length $\epsilon_{i}$. Then we have
$[c,b]\subseteq(a,b]=A\subseteq\sqcup_{i=1}^{\infty}A_{i}\subseteq\cup_{i=1}^{\infty}(a_{i},c_{i}).$
@@
Therefore, we conclude that $N(gJ_{1};I_{k})$ and $N(J;I_{k})$ differ by at most $62\eta N_{k}+$ 6. For the same reason, $N(gJ_{2};I_{k})$ and $N(S^{1}\setminus J;I_{k})$ differ by at most $62\eta N_{k}+$ 6. We now increase $k$ and observe
@@
We first take $\epsilon>0$ to be the constant as in Corollary 3.10. If a closed interval $[a,b]$ is $G$-contractible, then there exists $g\in G$ such that $\operatorname{\operatorname{diam}}(g[a,b])<\epsilon/3$. Because $g$ is continuous, we have $\operatorname{\operatorname{diam}}(g[\alpha,\beta])<\epsilon/2$ for $\alpha,\beta\in S^{1}\setminus[a,b]$ close enough to $a$ and $b$, respectively. Then Corollary 3.10 tells us that $g[\alpha,\beta]$ is $G$-contractible, i.e., there exists a sequence $\{g_{n}\}_{n>0}$ in $G$ such that $\lim_{n}\operatorname{\operatorname{diam}}(g_{n}\cdot g[\alpha,\beta])=0$. Now $\{g_{n}\cdot g\}_{n>0}$ is also a sequence in $G$ and contracts $[\alpha,\beta]$ arbitrarily small. Hence $[\alpha,\beta]$ is $G$-contractible. ∎
@@
Clearly $\{sg^{-1}s^{\prime}:s\in S_{1},s^{\prime}\in S_{i}\}\cup S_{i}$ is a finite subset of $G^{-1}$ for each $i$. Hence, a counterexample for $n$ leads to a counterexample for $n-1$. Since we have ruled out such a counterexample for $n=1$, the claim follows from the induction.
@@
Let $G$ be a subsemigroup of $\operatorname{Homeo}(S^{1})$ whose every $G$-orbit is infinite, and suppose that there exists a $G$-repeller $p$. We then claim that $G$ is proximal, i.e., $\inf_{g\in G}d(gx,gy)=0$ for each $x,y\in S^{1}$. First, when $x,y\neq p$ this is clear from the definition of $G$-repellers. If $x=p$, then we can pick some $h\in G$ such that $x\notin\{hx,hy\}$ using Lemma 3.13. It is easy to check that $h^{-1}x$ is also a $G$-repeller, which is distinct from $x$ and $y$. It is then clear that $\inf_{g\in G}d(gx,gy)=0$.
@@
Given $x\in S^{1}$, we take $z$ such that $Firm(x)=[x,z]$. Then by the assumption, $Firm(z)$ is not contained in $[z,x]$; there exists $a\in(x,z)$ such that $[z,a]$ is $G$-contractible. (Think of $a$ as a point “just right to $x$”). By Corollary 3.11, there exists $z^{\prime}\in(a,z)$ such that $[z^{\prime},a]$ is also $G$-contractible. (Think of $z^{\prime}$ as a point “just left to $z$”.) We now take $z^{\prime\prime}\in(z^{\prime},z)$. Then $z^{\prime\prime}\in S^{1}\setminus[z,a]\subseteq(x,z)$. Hence, $[x,z^{\prime\prime}]$ is $G$-contractible. This means $[x^{\prime},z^{\prime\prime}]$ is $G$-contractible for some $x^{\prime}\in(z^{\prime\prime},x)\subseteq(z,x)$ (Again, $x^{\prime}$ is “just left to $x$”).
To show that $U:=(x^{\prime},a)$ satisfies the desired property, let $I=(c,d)$ be an interval contained in $U$. Every $q\in S^{1}\setminus I$ either belongs to $[d,z^{\prime\prime}]$ or $[z^{\prime\prime},c]$. In the former case, $[p,q]\subseteq[x^{\prime},z^{\prime\prime}]$ is $G$-contractible for every $p\in I$. In the latter case, $[q,p]\subseteq[z^{\prime},a]$ is $G$-contractible for each $p\in I$. ∎
@@
Pick an arbitrary $x\in S^{1}$ and consider an increasing sequence of intervals $I_{n}=[x,z_{n}]$ that fills up the interior of $Firm(x)$. That means, writing $Firm(x)$ as $[x,z]$, we require that $[x,z_{n}]\subsetneq[x,z]$ and $\lim_{n}z_{n}=z$. Note that $z\notin[x,z_{n}]$. Each $I_{n}$ is $G$-contractible and there exists $g_{n}\in G$ such that $g_{n}I_{n}$ is $1/n$-short. By taking a subsequence, we may assume that $g_{n}I_{n}$ converges to a point $y$.
@@
In the former case, $[x,z]=Firm(x)$ becomes $G$-contractible, which is a contradiction. Hence, only the latter can be true: $[z,z_{n}]$ is $G$-contractible. Because this holds for arbitrary $n$ and because $[z,z_{n}]$ exhausts $S^{1}\setminus\{z\}$, we conclude that $z$ is a $G$-repeller. This contradicts the assumption. ∎
@@
Let $G$ be a subsemigroup of $\operatorname{Homeo}(S^{1})$. First note that a $G$-fixed point is by definition a singleton $G$-orbit. Next, if there exists a finite $G$-orbit, i.e., #$\#Gx<+\infty$ for some $x\in S^{1}$, then the uniform measure on $Gx$ becomes $G$-invariant. Indeed, for any $g\in G$, we have $g\cdot Gx\subseteq Gx$ by the semigroup property of $G$. Since $Gx$ is a finite set, this forces that $gGx=Gx$. In other words, each $g\in G$ permutes points in $Gx$ and preserves the uniform measure on $Gx$. We conclude the implication (3) $\Rightarrow$ (2) $\Rightarrow$ (1) in the statement.
@@
Next, given the objects for $n-1$, we define the ones for $n$. First, since $V_{n-1}$ is $\epsilon$-shot, it is $G$-contractible; there exists $\{h_{k}\}_{k>0}\subseteq G$ such that $\operatorname{\operatorname{diam}}(h_{k}V_{n-1})\rightarrow 0$. Up to subsequence, we may assume that $h_{k}V_{n-1}$ converges to a point $x_{n}$. Up to a further subsequence, we may assume that $\{h_{k}S_{n-1}\}_{k>0}$, a sequence of finite sets of cardinality #$\#S_{n-1}$, converges to a finite set $S^{\prime}$. We then define $S_{n}:=S^{\prime}\cup\{x_{n}\}$. For each $k$, we can take large enough $i(k)$ such that
@@
We have not yet put restrictions on the size of $S_{n}$, but by modifying the choice of $\{g_{k;n}\}_{k>0}$ (while keeping $K_{n}:=W_{0}\sqcup\ldots\sqcup W_{n-1}$) we can make #$\#S_{n}\leq N$. To see this, suppose that #$\#S_{n}>N$. Then one of $I_{1},\ldots,I_{N}$ contains more than 2 points of $S_{n}$. Without loss of generality, suppose that $I_{1}$ does so. Because $I_{1}$ has length $\epsilon$ and is $G$-contractible, there exist $\{h_{k}\}_{k>0}$ in $G$ such that $\operatorname{\operatorname{diam}}(h_{k}I_{1})\rightarrow 0$. Up to subsequence, we may assume that $h_{k}I_{1}$ converges to a point $x$. Moreover, $\{h_{k}(S_{n}\setminus I_{1})\}_{k>0}$ is a sequence of sets of cardinality at most #$\#S_{n}-2$. Up to subsequence, we may assume that $h_{k}(S_{n}\setminus I_{1})$ converges to a finite set $F$ of cardinality at most #$\#S_{n}-2$. We now set $S_{n}^{\prime}:=F\cup\{x\}$; then $h_{k}S_{n}\rightarrow S_{n}^{\prime}$ holds. For each $k$, we choose large enough $i(k)$ such that
@@
Then $\{h_{k}g_{i(k);n}\}_{k>0}$ is a sequence in $G$ such that $h_{k}g_{i(k);n}(K_{n})$ converges to $S_{n}^{\prime}$. Note that #$\#S_{n}^{\prime}\leq\#S_{n}-1$. By applying this procedure inductively, we can make #$\#S_{n}$ smaller than or equal to $N$. Hence, we may assume that #$\#S_{n}\leq N$ for each $n$.
We thus have a sequence $(K_{n})_{n>0}$ of elementary sets, a sequence $(S_{n})_{n>0}$ of sets with cardinality at most $N$, and a sequence ($(\{g_{k;n}\}_{k>0})_{n>0}$ of sequences in $G$ that satisfy the following for each $n$:
@@
Now, by replacing $(K_{n},S_{n},\{g_{k;n}\}_{k>0})_{n>0}$ with a suitable subsequence, we may assume that $S_{n}$ converges to a finite set $S$. (This is where the uniform bound on #$\#S_{n}$ is needed.) We then pick small enough $\eta>0$ and consider the $\eta$-neighborhood $N_{\eta}(S)$ of $S$. Then $S_{n}\subseteq N_{\eta/2}(S)$ for suitably large $n$, and $g_{k(n);n}(K_{n})\subseteq N_{\eta/2}(S_{n})\subseteq N_{\eta}(S)$ for suitably large $k(n)$. Hence, $L_{n}:=g_{k(n);n}^{-1}(S^{1}\setminus N_{\eta}(S))$ becomes a set of Lebesgue measure at most $1-\operatorname{Leb}(K_{n})$. In fact, this set is a union of #$\#S$ closed intervals, each with length at most $1-\operatorname{Leb}(K_{n})$. Let $C_{n}$ be the set of centers of these intervals. Up to subsequence, we may assume that $C_{n}$ converges to a finite subset $C$ as $n$ tends to infinity. Because $L_{n}$ lies in the $(1-\operatorname{Leb}(K_{n}))$-neighborhood of $C_{n}$, $L_{n}\subseteq N_{\eta}(C)$ for suitably large $n$.

### v2

@@ Abstract.
We adapt Gouëzel’s pivoting technique to the circle homeomorphism group. As an application, we give different proofs of Gilabert Vio’s probabilistic Tits alternative and Malicet’s exponential synchronization.
@@ 1. Introduction
$\operatorname{\mathbb{P}}\left(\textrm{$Z_{n}$ and $Z_{n}^{\prime}$ comprise a ping-pong pair}\right)\geq 1-q^{n}$
@@
$\lim_{N\rightarrow\infty}\frac{1}{N}\#\{0\leq n\leq N\,|\,\textrm{$Z_{n}$ and $Z_{n}^{\prime}$ comprise a ping-pong pair}\}=1.$
@@
(1.1) $\operatorname{\mathbb{P}}\left(\textrm{$Z_{n}$ and $Z_{n}^{\prime}$ comprise a ping-pong pair}\right)\geq 1-\frac{1}{\kappa}e^{-\kappa n}$
@@
In the first case, $f_{i}$’s simultaneously preserve a metric on $S^{1}$ and distinct points are kept distant. In the second case, a global synchronization cannot be expected but a local synchronization can happen. In the third case, synchronization happens almost surely. An independent work was done by V. A. Kleptsyn and M. B. Nalskii under the additional assumption that $f_{1},\ldots,f_{m}$ generate a hyperbolic map [KN04].
Synchronization in random dynamical systems was also studied by Bertrand Deroin, Victor Kleptsyn and Andrés Navas [DKN07, Proposition 5.2], where the authors gave an exponential upper bound on the probability of non-synchronization. The authors then related synchronization in $\mu$-random walk with properties of the $\mu$-stationary measure such as non-atomness and uniqueness. See also [Nav11] for further context and overview.
Synchronization was promoted into exponential synchronization by Dominique Malicet in the following form. Note that the minimality assumption is lifted.
@@
We strengthen this result by providing an exponential bound for exponential synchronization:
@@
[…]
@@
[…]
@@
[…]
@@ Acknowledgments
The author thanks Hyungryul Baik and Sang-hyun Kim for their valuable comments on the Korean version of this paper. The author thanks Dominque Malicet and Martín Gilabert Vio for explaining the basics of circle homeomorphisms and clarifying the author’s many confusions. Malicet and Gilabert Vio encouraged the author to consider non-degenerate random walks on subsemigroups (which are not necessarily subgroups) of $\operatorname{Homeo}(S^{1})$, which led to the current statements. The author is grateful to their suggestions. The author also thanks Gilabert Vio for many corrections and suggestions for the first draft of this paper. Lastly, the author is grateful to Étienne Ghys, Victor Kleptsyn and Andrés Navas for their comments on the first draft of this paper and reference suggestions.
The author was partially supported by Mid-Career Researcher Program (RS-2023-00278510) through the National Research Foundation funded by the government of Korea. The first draft of the paper was written while the author was visiting the HCMC of KIAS.
@@ 2. Preliminaries
Given a subset $A\subseteq S^{1}$, we denote by $\zeta(A)$ the number of connected components of $A$ (possibly $\infty$). Hence, $\zeta(A)=1$ precisely when $A$ is connected, i.e., is a point, interval or $S^{1}$.
@@ 3. Weak Tits alternative for semigroups
    $G$ acts on $S^{1}$ proximally and every $G$-orbit is infinite, i.e., $Gx=+\infty$ for each $x\in S^{1}$.
@@
[…]
@@
$\left(\frac{N(A;I_{n})}{N(S^{1};I_{n})}\right)_{n>0}=\left(\frac{N(A;I_{1})}{N(S^{1};I_{1})},\,\frac{N(A;I_{2})}{N(S^{1};I_{2})},\,\frac{N(A;I_{3})}{N(S^{1};I_{3})},\ldots\right)$
is well-defined thanks to Claim 3.6 and is bounded between 0 and 1. Hence, up to replacing $(I_{n})_{n>0}$ with its subsequence, $(N(A;I_{n})/N(S^{1};I_{n}))$ converges. Since $\mathcal{E}$ is countable, we can subsequently take convergent subsequences for each $A\in\mathcal{E}$. As a result, for a suitable interval neighborhood basis $(I_{n})_{n>0}$ of $x$, we can guarantee that
$\lim_{n\rightarrow+\infty}\frac{N(A;I_{n})}{N(S^{1};I_{n})}=:\mu(A)\,\,\textrm{exists for each $A\in\mathcal{E}$.}$
@@
Meanwhile, we observed in Claim 3.5 that $N(S^{1};I_{n})$ grows indefinitely as $n$ tends to infinity. Hence, we conclude
$0\leq\big{|}\mu(A_{1})+\mu(A_{2})-\mu(A)\big{|}=\lim_{n\rightarrow+\infty}\left|\frac{N(A_{1};I_{n})+N(A_{2};I_{n})-N(A;I_{n})}{N(S^{1};I_{n})}\right|\leq\lim_{n\rightarrow+\infty}\left|\frac{2}{N(S^{1};I_{n})}\right|=0.$
@@
Each $A_{i}$ is of the form $(a_{i},b_{i}]$. We now take $c_{i}\in S^{1}$ such that $b_{i},c_{i}]$ has length $\epsilon_{i}$. Then we have
$c,b]\subseteq(a,b]=A\subseteq\sqcup_{i=1}^{\infty}A_{i}\subseteq\cup_{i=1}^{\infty}(a_{i},c_{i}).$
@@
Therefore, we conclude that $N(gJ_{1};I_{k})$ and $N(J;I_{k})$ differ by at most $2\eta N_{k}+6$. For the same reason, $N(gJ_{2};I_{k})$ and $N(S^{1}\setminus J;I_{k})$ differ by at most $2\eta N_{k}+6$. We now increase $k$ and observe
@@
We first take $\epsilon>0$ to be the constant as in Corollary 3.10. If a closed interval $[a,b]$ is $G$-contractible, then there exists $g\in G$ such that $\operatorname{\operatorname{diam}}(g[a,b])<\epsilon/3$. Because $g$ is continuous, we have $\operatorname{\operatorname{diam}}(g[\alpha,\beta])<\epsilon/2$ for $\alpha,\beta\in S^{1}\setminus[a,b]$ close enough to $a$ and $b$, respectively. Then Corollary 3.10 tells us that $g[\alpha,\beta]$ is $G$-contractible, i.e., there exists a sequence $\{g_{n}\}_{n>0}$ in $G$ such that $\lim_{n}\operatorname{\operatorname{diam}}(g_{n}\cdot g[\alpha,\beta])=0$. Now $\{g_{n}\cdot g\}_{n>0}$ is also a sequence in $G$ and contracts $\alpha,\beta]$ arbitrarily small. Hence $\alpha,\beta]$ is $G$-contractible. ∎
@@
Clearly $\{sg^{-1}s^{\prime}:s\in S_{1},s^{\prime}\in S_{i}\}\cup S_{i}$ is a finite subset of $G^{-1}$ for each $i$. Hence, a counterexample for $n$ leads to a counterexample for n1n-1. Since we have ruled out such a counterexample for $n=1$, the claim follows from the induction.
@@
Let $G$ be a subsemigroup of $\operatorname{Homeo}(S^{1})$ whose every $G$-orbit is infinite, and suppose that there exists a $G$-repeller pp. We then claim that $G$ is proximal, i.e., $\inf_{g\in G}d(gx,gy)=0$ for each $x,y\in S^{1}$. First, when $x,y\neq p$ this is clear from the definition of $G$-repellers. If $x=p$, then we can pick some $h\in G$ such that $x\notin\{hx,hy\}$ using Lemma 3.13. It is easy to check that $h^{-1}x$ is also a $G$-repeller, which is distinct from $x$ and $y$. It is then clear that $\inf_{g\in G}d(gx,gy)=0$.
@@
Given $x\in S^{1}$, we take $z$ such that $Firm(x)=[x,z]$. Then by the assumption, $Firm(z)$ is not contained in $[z,x]$; there exists $a\in(x,z)$ such that $[z,a]$ is $G$-contractible. (Think of $a$ as a point “just right to xx”). By Corollary 3.11, there exists $z^{\prime}\in(a,z)$ such that $z^{\prime},a]$ is also $G$-contractible. (Think of $z^{\prime}$ as a point “just left to zz”.) We now take $z^{\prime\prime}\in(z^{\prime},z)$. Then $z^{\prime\prime}\in S^{1}\setminus[z,a]\subseteq(x,z)$. Hence, $x,z^{\prime\prime}]$ is $G$-contractible. This means $x^{\prime},z^{\prime\prime}]$ is $G$-contractible for some $x^{\prime}\in(z^{\prime\prime},x)\subseteq(z,x)$ (Again, $x^{\prime}$ is “just left to xx”).
To show that $U:=(x^{\prime},a)$ satisfies the desired property, let $I=(c,d)$ be an interval contained in $U$. Every $q\in S^{1}\setminus I$ either belongs to $d,z^{\prime\prime}]$ or $z^{\prime\prime},c]$. In the former case, $p,q]\subseteq[x^{\prime},z^{\prime\prime}]$ is $G$-contractible for every $p\in I$. In the latter case, $q,p]\subseteq[z^{\prime},a]$ is $G$-contractible for each $p\in I$. ∎
@@
Pick an arbitrary $x\in S^{1}$ and consider an increasing sequence of intervals $I_{n}=[x,z_{n}]$ that fills up the interior of $Firm(x)$. That means, writing $Firm(x)$ as $[x,z]$, we require that $x,z_{n}]\subsetneq[x,z]$ and $\lim_{n}z_{n}=z$. Note that $z\notin[x,z_{n}]$. Each $I_{n}$ is $G$-contractible and there exists $g_{n}\in G$ such that $g_{n}I_{n}$ is $1/n$-short. By taking a subsequence, we may assume that $g_{n}I_{n}$ converges to a point $y$.
@@
In the former case, $[x,z]=Firm(x)$ becomes $G$-contractible, which is a contradiction. Hence, only the latter can be true: $z,z_{n}]$ is $G$-contractible. Because this holds for arbitrary $n$ and because $z,z_{n}]$ exhausts $S^{1}\setminus\{z\}$, we conclude that $z$ is a $G$-repeller. This contradicts the assumption. ∎
@@
Let $G$ be a subsemigroup of $\operatorname{Homeo}(S^{1})$. First note that a $G$-fixed point is by definition a singleton $G$-orbit. Next, if there exists a finite $G$-orbit, i.e., $Gx<+\infty$ for some $x\in S^{1}$, then the uniform measure on $Gx$ becomes $G$-invariant. Indeed, for any $g\in G$, we have $g\cdot Gx\subseteq Gx$ by the semigroup property of $G$. Since $Gx$ is a finite set, this forces that $gGx=Gx$. In other words, each $g\in G$ permutes points in $Gx$ and preserves the uniform measure on $Gx$. We conclude the implication (3) $\Rightarrow$ (2) $\Rightarrow$ (1) in the statement.
@@
Next, given the objects for n1n-1, we define the ones for $n$. First, since $V_{n-1}$ is $\epsilon$-shot, it is $G$-contractible; there exists $\{h_{k}\}_{k>0}\subseteq G$ such that $\operatorname{\operatorname{diam}}(h_{k}V_{n-1})\rightarrow 0$. Up to subsequence, we may assume that $h_{k}V_{n-1}$ converges to a point $x_{n}$. Up to a further subsequence, we may assume that $\{h_{k}S_{n-1}\}_{k>0}$, a sequence of finite sets of cardinality $S_{n-1}$, converges to a finite set $S^{\prime}$. We then define $S_{n}:=S^{\prime}\cup\{x_{n}\}$. For each $k$, we can take large enough $i(k)$ such that
@@
We have not yet put restrictions on the size of $S_{n}$, but by modifying the choice of $\{g_{k;n}\}_{k>0}$ (while keeping $K_{n}:=W_{0}\sqcup\ldots\sqcup W_{n-1}$) we can make $S_{n}\leq N$. To see this, suppose that $S_{n}>N$. Then one of $I_{1},\ldots,I_{N}$ contains more than 2 points of $S_{n}$. Without loss of generality, suppose that $I_{1}$ does so. Because $I_{1}$ has length $\epsilon$ and is $G$-contractible, there exist $\{h_{k}\}_{k>0}$ in $G$ such that $\operatorname{\operatorname{diam}}(h_{k}I_{1})\rightarrow 0$. Up to subsequence, we may assume that $h_{k}I_{1}$ converges to a point $x$. Moreover, $\{h_{k}(S_{n}\setminus I_{1})\}_{k>0}$ is a sequence of sets of cardinality at most $S_{n}-2$. Up to subsequence, we may assume that $h_{k}(S_{n}\setminus I_{1})$ converges to a finite set $F$ of cardinality at most $S_{n}-2$. We now set $S_{n}^{\prime}:=F\cup\{x\}$; then $h_{k}S_{n}\rightarrow S_{n}^{\prime}$ holds. For each $k$, we choose large enough $i(k)$ such that
@@
Then $\{h_{k}g_{i(k);n}\}_{k>0}$ is a sequence in $G$ such that $h_{k}g_{i(k);n}(K_{n})$ converges to $S_{n}^{\prime}$. Note that $S_{n}^{\prime}\leq\#S_{n}-1$. By applying this procedure inductively, we can make $S_{n}$ smaller than or equal to $N$. Hence, we may assume that $S_{n}\leq N$ for each $n$.
We thus have a sequence $(K_{n})_{n>0}$ of elementary sets, a sequence $(S_{n})_{n>0}$ of sets with cardinality at most $N$, and a sequence $(\{g_{k;n}\}_{k>0})_{n>0}$ of sequences in $G$ that satisfy the following for each $n$:
@@
Now, by replacing $(K_{n},S_{n},\{g_{k;n}\}_{k>0})_{n>0}$ with a suitable subsequence, we may assume that $S_{n}$ converges to a finite set $S$. (This is where the uniform bound on $S_{n}$ is needed.) We then pick small enough $\eta>0$ and consider the $\eta$-neighborhood $N_{\eta}(S)$ of $S$. Then $S_{n}\subseteq N_{\eta/2}(S)$ for suitably large $n$, and $g_{k(n);n}(K_{n})\subseteq N_{\eta/2}(S_{n})\subseteq N_{\eta}(S)$ for suitably large $k(n)$. Hence, $L_{n}:=g_{k(n);n}^{-1}(S^{1}\setminus N_{\eta}(S))$ becomes a set of Lebesgue measure at most $1-\operatorname{Leb}(K_{n})$. In fact, this set is a union of #S\#S closed intervals, each with length at most $1-\operatorname{Leb}(K_{n})$. Let $C_{n}$ be the set of centers of these intervals. Up to subsequence, we may assume that $C_{n}$ converges to a finite subset $C$ as $n$ tends to infinity. Because $L_{n}$ lies in the $(1-\operatorname{Leb}(K_{n}))$-neighborhood of $C_{n}$, $L_{n}\subseteq N_{\eta}(C)$ for suitably large $n$.